# Permutation twisted cohomology, remixed

### v2

@@ 6. Injectivity of the comparison map
By construction $\hat{\Psi}^{({}^{g}H)}(f^{\prime\prime})$ is an isomorphism for all $g\in H_{i+1}\setminus H_{i}$, therefore $^{g}(f^{\prime\prime})\not\in\mathfrak{q}$. But since $\mathfrak{p},\mathfrak{q}$ satisfy $(\ast)$, it follows that $f^{\prime\prime}\in\mathfrak{q}$ as well. Since each $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{q}$, cancellation implies $f\in\mathfrak{q}$ as well. Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$, and the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, as desired. ∎
As an immediate corollary, we obtain a strengthening of Proposition 5.15.
@@
Let $H$ be any subgroup of $G$. Then we have a commutative square
@@
In particular, the diagonal is injective, and the product
$\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective.
@@
In particular, $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points (maximal primes)
$\operatorname{comp}_{G}(\mathfrak{m}_{H})=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
@@
$\operatorname{comp}_{\mathscr{L}_{G}(H)}:\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is noetherian as well, hence [Bal10a, Theorem 7.3] implies $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a continuous bijection. It remains to show this map is a homeomorphism; we show it is closed. Let $\mathscr{L}_{G}(H)_{\langle k\rangle}$ denote the unitization of $\mathscr{L}_{G}(H)$, i.e. the tt-subcategory compactly generated by the tensor unit. Of course, the unitization has the same cohomology ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$. Then [San25, Corollary 10.2] asserts that the inclusion $\mathscr{L}_{G}(H)_{\langle k\rangle}\to\mathscr{L}_{G}(H)$ induces a homeomorphism $\operatorname{Spc}(\mathscr{L}_{G}(H)_{\langle k\rangle})\cong\operatorname{Spc}(\mathscr{L}_{G}(H))$. The comparison map factors through $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}$ as follows:
$\operatorname{Spc}(\mathscr{L}_{G}(H))\to\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
Now, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is (by definition) generated by its tensor unit. By noetherianity of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is therefore End-finite in the sense of [Lau23, Definition 2.6], hence [Lau23, Proposition 2.7] asserts the induced map on spectra $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism, thus $\operatorname{comp}_{\mathscr{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism as well, as desired.
Therefore, $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. It is easy to check the closed points $\mathfrak{m}_{H}$ are mapped to the corresponding maximal primes in the image as described in the theorem statement. ∎

### v3

@@ 6. Injectivity of the comparison map
By construction $\hat{\Psi}^{({}^{g}H)}(f^{\prime\prime})$ is an isomorphism for all $g\in H_{i+1}\setminus H_{i}$, therefore $^{g}(f^{\prime\prime})\not\in\mathfrak{q}$. But since $(\mathfrak{p},\mathfrak{q})$ satisfies $(\ast)$, it must be the case that $f^{\prime\prime}\in\mathfrak{q}$. Since each $^{g}(\iota^{H}_{C})\in\mathscr{P}(H,0)\subseteq\mathfrak{q}$, cancellation or homogeneity of $\mathfrak{p}$ implies $f\in\mathfrak{q}$ as well. Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$. Putting everything together, we conclude the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective. ∎
Note that the restriction $\rho_{N_{G}(H)}^{G}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ need not be injective, as fusion phenomena will occur. As an immediate corollary, we strengthen 5.10 and 5.11.
@@
Let $H\leq G$ be any subgroup. Then the commutative square
@@
satisfies that the diagonal is injective. Moreover,
$\hat{\psi}^{h}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective with dense image.
Remark 6.5.
The comparison map is an injective map of spectral spaces, therefore Stone duality implies the pullback
$\Omega(\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)))\to\Omega(\operatorname{Spc}(\mathscr{K}(G))),\,U\mapsto\operatorname{comp}_{G}^{-1}(U)$
is an epimorphism in CFrm, where $\Omega(X)$ denotes the coherent frame of open subsets of a spectral topological space $X$. However, epimorphisms in the category CFrm of bounded distributive lattices need not be surjective. This is in contrast to the category of posets, where epimorphisms are precisely the surjective maps.
In fact, if the induced map on opens is surjective, then by definition, the submonoid of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ of effective endotrivials is by definition an ample family [GAL25, Definition 5.4], and by [GAL25, Proposition 5.8], the comparison map would be an open immersion. Therefore, in order to completely deduce the topology of $\operatorname{Spc}(\mathscr{K}(G))$, more work is required.
On the other hand, by [DST19, Theorem 5.2.2], we have the comparison map induces a surjective map
$\mathcal{K}(\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)))\to\mathcal{K}(\operatorname{Spc}(\mathscr{K}(G))),$
where $\mathcal{K}(X)$ denotes the poset of clopen subsets of $X$.
If we know twisted cohomology is Noetherian, we further can show that the comparison map restricts to a homeomorphism upon localization to the opens $U(H)$, and therefore is an open immersion, extending Balmer–Gallauer’s [BG25, Theorem 15.3]. The proof of this is quite short, thanks to machinery recently developed by Sanders [SAN25] regarding fully faithful tt-functors.
@@
In particular, in this case $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points
$\operatorname{comp}_{G}(\mathscr{M}(H))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
@@
$\operatorname{comp}_{\mathscr{L}_{G}(H)}\colon\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is Noetherian as well. Now, a theorem of Sanders [SAN25, Theorem 10.8] immediately $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a homeomorphism, as desired. Thus, globally $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. It is easy to check the closed points $\mathscr{M}(H)$ are mapped to the corresponding maximal primes in the image as described in the theorem statement. ∎